OpenStudy (anonymous):

(a+b) + (c-d) - (a-b) - (c+d) + (a-b) simplify ff by removing brackets and collecting terms. Please help

OpenStudy (binary3i):

(a+b) + (c-d) - (a-b) - (c+d) + (a-b) = a+b+c-d-a+b-c-d+a-b= b-d-d+a=a+b-2d

OpenStudy (anonymous):

thanks... @binary3i

OpenStudy (binary3i):

welcome